# Information Science Principles of Machine Learning: A Causal Chain Meta-Framework Based on Formalized Information Mapping

### v9

@@ 5.2 Ethical Safety Assurance Experiment
Given the nature of the bivariate root function regression task, the model’s predicted outputs should never be negative. This requirement is analogous to Definition 10, which ensures that model outputs do not violate the ethical constraints . Therefore, using the same training dataset as in the previous prediction error experiments, we set the learning rate to and trained the model for 165 steps. For testing, we generated a dataset of 10,000 samples with drawn uniformly from within a circle of radius 0.5 centered at (-1, -1). The criterion for judging ethical safety violations was whether the model’s predicted output was negative. This setup was used to test the feasibility of Theorem 5 and to evaluate the additional computational overhead introduced by ethical safety checking.
@@
Trial Computation Computation Computation Number of Number of Correct Correct
Number Time Without Time With Time Ratio Ethically Unethical Prediction Prompt
Ethical Ethical Safe Warning Rate for Rate for
Safety Check Safety Check Samples Samples Safe Unethical
(seconds) (seconds) Outputs Outputs
@@ 5.3.1 Mathematical Operation-Based Explanation
$Y=\begin{bmatrix}y_{1}\\ y_{2}\end{bmatrix}=\text{Identity}(W_{2}\text{ReLU}(W_{1}X+b_{1})+b_{2})\approx\begin{bmatrix}-0.040\\ 0.277\end{bmatrix}$ (5.3)
@@
These formulas provide the designer’s mathematical representation of the model’s computational process. They are equivalent to the formal forward propagation rules (A.21) and (A.22) in the appendix. According to the OIT perspective, these equations describe the ontological state of the model, its outputs, and the information they convey. The model, along with its input and output mechanisms, serves as the carrier of this information and realizes the computational process. Therefore, using these formulas to explain why the model produces negative outputs satisfies the requirements of Definition 9 regarding model interpretability and meets the needs of researchers who seek to understand model behavior from a mathematical standpoint.
@@ 5.3.2 Visualization-Based Explanation of the Computational Process
Depicting the model’s computational process visually can make explanations of model behavior much more intuitive. To this end, for model M after 165 training steps, we illustrate its complete structure—covering the inputs, internal neural network calculations and data flow, and outputs—in Figure 3.
@@
Compared with equations (5.1)(5.3), Figure 3 not only comprehensively reflects the input, weight matrices, bias vectors, neuron activation functions, and the specific data and formulas for outputs, but also highlights the interactions between the input/output devices and the model. The figure offers a more direct view of the internal data processing pipeline, making it easier to examine and verify each step of the computation. Thus, the visualization provides a more accessible and concrete explanation of how model M processes inputs and generates results.
@@ 5.3.3 Partition-Based Explanation
It is evident that equations (5.1)(5.3) and Figure3 explain only the result of model M processing the specific sample $X=[-1.3535,-1.3535]^{T}$. They do not account for the phenomenon observed in the ethical safety assurance experiment, where each random selection of 5,000 samples from the test dataset yields over 19% negative predictions. To explain this, a broader statistical analysis across more data samples is needed.
Figure 4(a) provides a visual depiction of the decision boundary after 165 training steps, showing where model M predicts negative outputs. Notably, there is a boundary line approximately passing through points (-2.0, 0.4) and (-0.4, -2.0). For samples on the upper right side of this diagonal, model M consistently predicts non-negative values; for those on the lower left side, the model consistently outputs negative values. The test region in the figure, representing the test dataset, falls precisely on the lower left side of this boundary line, which explains why a random selection of 5,000 samples from the test set will always include a substantial portion that triggers negative predictions. This visual and statistical partitioning thus provides an intuitive explanation for the observed frequency of negative outputs in the ethical safety experiment.
Figure 4(a) presents an overall view of numerous inputs processed and predicted by model M. This visualization also represents an ontological state of M as it handles a wider range of input information—the real processing and prediction procedure of M is a direct result of this ontological mapping. Thus, according to Definition 9, Figure 4(a) constitutes a valid explanation for the distribution of prediction polarities by M.
@@ 5.3.4 Explanation Based on Cumulative Training Effects
The previous three explanations all focus on the behavior of M after 165 training steps. However, model training is a continuous process of cumulative improvement. Examining only the model’s instantaneous performance at a particular training step cannot fully explain the causes of abnormal outputs. To address this, we further trained M for 400 iterations on the training dataset. Figure 4(b) shows the predictions for both $(1,1)$ and (1.3535,1.3535)(-1.3535,-1.3535) after each training step.
@@
It is clear that even for $(1,1)$ — a central sample from the training set — the predicted $y_{2}$ by M remains negative during roughly the first 50 training steps. Only after more than 50 steps does the prediction move away from negative values. For the test sample (1.3535,1.3535)(-1.3535,-1.3535), it takes more than about 180 training steps before the prediction begins to consistently avoid negative values. This cumulative effect is entirely due to the optimization of M’s parameters as training progresses.
@@
In fact, by applying the forward propagation rules (A.21) and (A.22) from the appendix—substituting the model parameters obtained after each training step—one can, through a recursive process, provide a rigorous and step-by-step explanation for M’s predictions on specific input samples after any given amount of training.
6 Conclusion and Outlook
This paper has established the "Meta-Framework for Machine Learning Theory (MLT-MF)" and, through formal information modeling and the sextuple definition, has unified the characterization of information flow and causal mechanisms across the entire machine learning lifecycle. The core contributions are as follows:
  • Theoretical Integration: For the first time, data initialization, training optimization, and knowledge transfer are incorporated into a unified framework, overcoming the fragmented modeling of temporal causal chains in traditional research;
  • Interpretability: By explicitly representing ontological and carrier states via higher-order logical systems and mapping model decisions back to the formal semantic space through information recoverability, logical transparency is ensured and a rigorous mathematical foundation for interpretability is provided;
  • Ethical Safety: By embedding ethical constraint predicates, the compliance of model outputs can be formally verified, offering criteria and methodological guidance for the design of trustworthy AI systems;
  • Model Performance: Based on state coverage and KL divergence, an upper bound formula for TVD is derived, revealing the core principle that the “problem space must be contained within the model knowledge space,” thus providing explicit tools for quantitative analysis in data collection and model optimization.
We also validated and demonstrated the correctness and practicality of the relevant theorems based on MLT-MF through experiments with a simple feedforward neural network model.
Despite these important advances, the current framework assumes ideal (noiseless) enabling mappings and primarily addresses model learning and processing logic in static scenarios. It does not yet address information fusion and conflict resolution across ontological spaces in multimodal or multi-agent systems. Future research should deepen the following directions:
  • Extend to complex dynamic scenarios such as online learning under noisy conditions, introducing noisy ontologies and incremental coverage mechanisms;
  • Investigate information fusion and conflict resolution across ontological spaces for multimodal and multi-agent systems, constructing distributed MLT-MF frameworks;
  • Promote interdisciplinary integration of MLT-MF with fields such as quantum computing and bioinformatics, exploring formal mappings of quantum states and interpretable modeling of gene regulatory networks;
  • Design automated toolchains to support conversion from formal rules to code, lowering the engineering threshold for MLT-MF adoption.
7 Acknowledgements
@@
During the writing and revision of this paper, I received many insightful comments from Associate Professor Wang Rui of the School of Computer Science at Shanghai Jiao Tong University. I also gained much inspiration and assistance from regular academic discussions with doctoral students Siyuan Qiu, Chun Li, Hu Xu, and Zeyan Li. I hereby express my sincere gratitude to them. Junxuan He from Beijing University of Posts and Telecommunications implemented all the experiments described in the paper according to the author’s ideas, producing complete model code and training/test datasets, providing indispensable and strong support for illustrating and validating the theorems related to MLT-MF. The author offers special thanks for his hard work, keen understanding, and efficient implementation.
8 Data and Code availability
@@
Appendix A Appendix: Model Instance Validation of MLT-MF
@@
A.1 A Simple Feedforward Neural Network Model
We choose a feedforward neural network M with an input layer containing two inputs, a hidden layer containing three neurons, and an output layer containing two outputs (Figure A1). Its simplicity helps explain the essence of MLT-MF in a clear and straightforward manner.
Figure A1: Schematic diagram of feedforward neural network model M

### v10

@@ 6.3 Ethical Safety Assurance Experiment
Given the nature of the bivariate root function regression task, the model’s predicted outputs should never be negative. This requirement is analogous to Definition 5, which ensures that model outputs do not violate the ethical constraints $Ec$. Therefore, using the same training dataset as in the previous prediction error experiments, we set the learning rate to $\eta=0.001$ and trained the model for 165 steps. For testing, we generated a dataset of 10,000 samples with $(x_{1},x_{2})$ drawn uniformly from within a circle of radius 0.5 centered at (-1, -1). The criterion for judging ethical safety violations was whether the model’s predicted output was negative. This setup was used to test the feasibility of Theorem 2 and to evaluate the additional computational overhead introduced by ethical safety checking.
@@
Trial Time w/o Time w/ Time Safe Warning Safe Warning
Number Check (s) Check (s) Ratio Samples Samples Rate Rate
@@ 6.4.1 Mathematical Operation-Based Explanation
$Y=\begin{bmatrix}y_{1}\\ y_{2}\end{bmatrix}=\text{Identity}\left(W_{2}\text{ReLU}\left(W_{1}X+b_{1}\right)+b_{2}\right)\approx\begin{bmatrix}-0.040\\ 0.277\end{bmatrix}$ (6.3)
@@
These formulas provide the designer’s mathematical representation of the model’s computational process. They are equivalent to the formal forward propagation rules (C.21) and (C.22) in Appendix C. According to the OIT perspective, these equations describe the ontological state of the model, its outputs, and the information they convey. The model, along with its input and output mechanisms, serves as the carrier of this information and realizes the computational process. Therefore, using these formulas to explain why the model produces negative outputs satisfies the requirements of Definition 4 regarding model interpretability and meets the needs of researchers who seek to understand model behavior from a mathematical standpoint.
@@ 6.4.2 Visualization-Based Explanation of the Computational Process
Depicting the model’s computational process visually can make explanations of model behavior much more intuitive. To this end, for model M after 165 training steps, we illustrate its complete structure—covering the inputs, internal neural network calculations and data flow, and outputs—in Fig. 5.
@@
Compared with equations (6.1)–(6.3), Fig. 5 not only comprehensively reflects the input, weight matrices, bias vectors, neuron activation functions, and the specific data and formulas for outputs, but also highlights the interactions between the input/output devices and the model. The figure offers a more direct view of the internal data processing pipeline, making it easier to examine and verify each step of the computation. Thus, the visualization provides a more accessible and concrete explanation of how model M processes inputs and generates results.
@@ 6.4.3 Partition-Based Explanation
It is evident that equations (6.1)–(6.3) and Fig. 5 explain only the result of model M processing the specific sample $X=[-1.3535,-1.3535]^{T}$. They do not account for the phenomenon observed in the ethical safety assurance experiment, where each random selection of 5,000 samples from the test dataset yields over 19% negative predictions. To explain this, a broader statistical analysis across more data samples is needed.
Fig. 6(a) provides a visual depiction of the decision boundary after 165 training steps, showing where model M predicts negative outputs. Notably, there is a boundary line approximately passing through points (-2.0, -0.4) and (-0.4, -2.0). For samples on the upper right side of this diagonal, model M consistently predicts non-negative values; for those on the lower left, the model consistently outputs negative values. The red region in the figure, representing the test dataset, falls precisely on the lower left side of this boundary line, which explains why a random selection of 5,000 samples from the test set will always include a substantial portion that triggers negative predictions. This visual and statistical partitioning thus provides an intuitive explanation for the observed frequency of negative outputs in the ethical safety experiment.
@@
Fig. 6(a) presents an overall view of numerous inputs processed and predicted by model M. This visualization also represents an ontological state of M as it handles a wide range of input information—the real processing and prediction procedure of M is a direct result of this ontological mapping. Thus, according to Definition 4, Figure 5(a) constitutes a valid explanation for the distribution of prediction polarities by M.
@@ 6.4.4 Explanation Based on Cumulative Training Effects
The previous three explanations all focus on the behavior of M after 165 training steps. However, model training is a continuous process of cumulative improvement. Examining only the model's instantaneous performance at a particular training step cannot fully explain the causes of abnormal outputs. To address this, we further trained M for 400 iterations on the training dataset. Fig. 6(b) shows the predictions for both (1, 1) and (-1.3535, -1.3535) after each training step.
@@
It is clear that even for (1, 1) — a central sample from the training set—the predicted $y_{2}$ by M remains negative during roughly the first 50 training steps. Only after more than 50 steps does the prediction move away from negative values. For the test sample (-1.3535, -1.3535), it takes more than about 180 training steps before the prediction begins to consistently avoid negative values. This cumulative effect is entirely due to the optimization of M’s parameters as training progresses.
@@
In fact, by applying the forward propagation rules (C.21) and (C.22) from the appendix—substituting the model parameters obtained after each training step—one can, through a recursive process, provide a rigorous and step-by-step explanation for M’s predictions on specific input samples after any given amount of training.
7 Discussion and Conclusion
This work, grounded in the Objective Information Theory (OIT), develops the MLT-MF framework, employing a formal logical system to integrate the complete causal chain of model construction, training, input processing, output, and online learning. Under a unified semantics, we propose formal definitions of model interpretability and ethical safety, derive two classes of explicit TVD upper bound theorems, and validate the framework through experiments with feedforward neural network models. The core contributions are as follows:
(1) Enabled Mapping for Information Transmission: We characterize the mathematical mapping and the dual physical-causal nature of information transfer from ontological states to carrier states through the concept of “enabled mapping.” This is further extended to chain transmission, mapping each engineering stage of model development to a segment of the information chain, such that model behavior, training, and inference are reduced to logically verifiable compositions of mappings.
(2) Unified Representation of Interpretability and Safety: By adopting higher-order formal logic as a unified representation of “state,” we propose necessary and sufficient conditions for interpretability—linking input, output, and model within the same formal system via logical closure—and construct maximal safe outputs for ethical safety based on hypergraph maximum independent sets.
(3) Integration of Statistical Error, Noise Structure, and Coverage Geometry: We integrate statistical learning error characterization with information noise structure and training coverage geometry, providing a closed-form TVD upper bound, and operationally decomposing it in terms of coverage fitting error, uncovered mass, KL divergence, and noise deletion/addition terms.
Therefore, this study addresses two longstanding gaps in machine learning theory: the lack of unified semantics and causal chain constraints across different stages, and the absence of verifiable mathematical definitions and constructive guarantees for ethical safety and interpretability. By unifying the “existence–transmission–realization” of information within formal logical explanations, we rigorously establish that model interpretability corresponds to information existence, thereby providing a logical-theoretical foundation for translating ethical safety from normative requirements into decidable and constructible guarantees. Using TVD as a fundamental measure of distributional alignment, and establishing correspondences with practical engineering concerns such as training coverage gaps, uncovered-region prediction strategies, and noise deletion/addition, this work bridges the gap between abstract generalization theory and data coverage/system implementation, thus connecting statistical learning, mathematical logic, information science, and systems engineering.
All conclusions and methods presented are amenable to engineering implementation:
  • Interpretability can be realized at multiple levels through unified formal systems, including mathematical formulas, dataflow diagrams, and parameter representations, which is especially important for high-risk industries.
  • Ethical safety is implemented via conflict detection and maximal independent set approximation of output propositions, enabling a deployable process for maximal safe outputs.
[…]
Of course, this work has limitations: the current formal system is most transparent for finite discrete sets, and further refinement is needed for continuous, time-varying, multimodal, and multi-agent scenarios, particularly regarding measurability and dynamic consistency. The maximal independent set for ethical safety is generally NP-hard, requiring the development of complexity-controlled approximation algorithms and alignment with business ethics standards. TVD upper bounds rely on conservative inequalities such as Pinsker’s; in typical scenarios, error estimates may be tightened using related theoretical advances. Physical causality in enabled mappings calls for interdisciplinary integration with various branches of physics, leveraging parameters such as energy consumption, latency, and uncertainty to advance from qualitative causality to quantitative verification. Future work may include large-scale empirical studies to verify the observability of TVD upper bounds and the benefits of strategy optimization, as well as further development of cross-modal and multi-agent information chain composition and conflict resolution methods, thereby enhancing the universality and practical value of MLT-MF in complex systems.
Acknowledgments and Disclosure of Funding
@@ 7 Discussion and Conclusion
Junxuan He from Beijing University of Posts and Telecommunications implemented all the experiments described in the paper according to the author’s ideas, producing complete model code and training/test datasets, providing indispensable and strong support for illustrating and validating the theorems related to MLT-MF. The author offers special thanks for his hard work, keen understanding, and efficient implementation. During the writing and revision of this paper, I received many insightful comments from Associate Professor Wang Rui of the School of Computer Science at Shanghai Jiao Tong University. I also gained much inspiration and assistance from regular academic discussions with doctoral students Siyuan Qiu, Chun Li, Hu Xu, and Zeyan Li. I hereby express my sincere gratitude to them.
Data and Code availability
@@
Appendix A Appendix A. Formal Representation of States
Definition A.1 (Higher-Order Formal Logical System): Suppose a formal system $\mathcal{L}$ is described using the following symbol set:
Variables $x_{1}$, $x_{2}$,…;
Individual constants $a_{1}$, $a_{2}$,…;
Functions $f_{1}^{1}$, $f_{2}^{1}$, … $f_{1}^{2}$, $f_{2}^{2}$, … $f_{1}^{3}$, $f_{2}^{3}$, …;
Parentheses $(,)$;
Predicates $A_{1}^{1}$, $A_{2}^{1}$, … $A_{1}^{2}$, $A_{2}^{2}$, … $A_{1}^{3}$, $A_{2}^{3}$, …;
Logical connectives “not” $\sim$, “implication” $\rightarrow$;
Quantifier “for all” $\forall$.
A term in $\mathcal{L}$ is defined as follows:
(1) Variables and individual constants are terms;
(2) If $f_{i}^{n}(n>0,i>0)$ is a function in $\mathcal{L}$, and $t_{1}$,…, $t_{n}$ are terms in $\mathcal{L}$ that do not contain $f_{i}^{n}$ itself, then $f_{i}^{n}(t_{1}$,…, $t_{n})$ is also a term in $\mathcal{L}$;
(3) If $A_{i}^{n}(n>0,i>0)$ is a predicate in $\mathcal{L}$, and $t_{1}$,…, $t_{n}$ are terms in $\mathcal{L}$ that do not contain $A_{i}^{n}$ itself, then $A_{i}^{n}(t_{1}$,…, $t_{n})$ is also a term in $\mathcal{L}$;
(4) The set of all terms in $\mathcal{L}$ is generated by (1), (2), and (3).
Based on this, we can define atomic formulas in $\mathcal{L}$:
If $A_{i}^{n}(n>0,i>0)$ is a predicate in $\mathcal{L}$ and $t_{1}$,…, $t_{n}$ are terms in $\mathcal{L}$ not containing $A_{i}^{n}$ itself, then $A_{i}^{n}(t_{1}$,…, $t_{n})$ is an atomic formula in $\mathcal{L}$.
We then define well-formed formulas (WFFs) in $\mathcal{L}$:
(5) Every atomic formula in $\mathcal{L}$ is a WFF;
[…]
(7) The set of all WFFs in $\mathcal{L}$ is generated by (5) and (6).
The higher-order formal logical system $\mathcal{L}$ defined here, based on the flexibility of predicate logic, can describe various relationships and constraints among states. Its recursive structure and higher-order quantification support the nested definition of complex data and behaviors, and it explicitly prohibits any predicate or function symbol from applying to itself—directly or indirectly—to avoid self-referential paradoxes.
For simplicity, $\mathcal{L}$ does not define connectives such as “and” $\land$, “or” $\vee$, “if and only if” $\leftrightarrow$, or the quantifier “exists” $\exists$, and other commonly used logical symbols. However, mathematical logic theory demonstrates that these symbols are entirely equivalent to appropriate combinations of existing symbols and can make many formulas more concise (Hamilton, 1988). Therefore, they are introduced as defined symbols in subsequent discussions.
Definition A.2 (Interpretation of a Formal System) An interpretation $E$ of a formal system $\mathcal{L}$ over a domain $D_{E}$ assigns specific values to constants, elements of the domain $D_{E}$ to variables, and turns function and predicate symbols into functions and predicates over that domain $D_{E}$, thereby yielding a set of WFFs each with a truth value.
Definition A.2 assigns concrete values and meaning to a set of formulas in a formal system $\mathcal{L}$ through interpretation, thereby providing a fundamental approach for formally describing the definitional, attributive, structural, relational, and evolutionary state properties of objects in both the subjective and objective worlds.
Postulate A.1 (Expression of Object States) The state set of an object collection over a specified time set is constituted by its attributes—such as properties, structures, values, and relations—as well as the evolution rules and logical combinations among these attributes, and satisfies the following conditions:
Lemma A.1 (Formal Expression of States) If the state set $S(X,T)$ of objects $X$ over times $T$ satisfies Postulate A.1, then it can be represented as an interpretation of a formal system $\mathcal{L}$ over the domain $X\times T$, and it is logically coherent.
Proof: Since the state set $S(X,T)$ of the object set $X$ over the time set $T$ satisfies Postulate A.1, we construct the corresponding formal system $\mathcal{L}$ as follows:
By (3), the functions and predicates in $\mathcal{L}$ represent the properties, forms, values, and relations of $X$ over $T$ as captured in $S(X,T)$.
By (4), consider two ternary predicates $A_{1}^{3}$ and $A_{2}^{3}$. Let $A_{1}^{3}(x_{1},t_{1},f_{1}^{2}\left(x_{1},t_{1}\right))$ and $A_{2}^{3}(x_{2},t_{2},f_{2}^{2}\left(x_{2},t_{2}\right))$ be two properties of $S(X,T)$ corresponding to objects $x_{1},x_{2}\in X$, times $t_{1},t_{2}\in T(t_{1}<t_{2})$, and binary functions $f_{1}\left(x_{1},t_{1}\right),f_{2}\left(x_{1},t_{1}\right)$. Suppose all $A_{1}^{3}(x_{1},t_{1},f_{1}^{2}\left(x_{1},t_{1}\right))$ satisfy the rule of evolving from $t_{1}$ to $t_{2}$ into some $A_{2}^{3}(x_{2},t_{2},f_{2}^{2}\left(x_{2},t_{2}\right))$. The formula $(\forall A_{1}^{3}\forall f_{1}^{2}\forall x_{1}A_{1}^{3}(x_{1},t_{1},f_{1}^{2}\left(x_{1},t_{1}\right))\rightarrow{\exists A_{2}^{3}\exists f_{2}^{2}\exists x_{2}A}_{2}^{3}(x_{2},t_{2},f_{2}^{2}\left(x_{2},t_{2}\right))$ is also a well-formed formula in $\mathcal{L}$, with respect to the interpretation of $x_{1},x_{2},t_{1},t_{2}$, and is able to express the corresponding rules in $S(X,T)$. For other rules in $S(X,T)$, the related elements in the domain $X\times T$ can similarly be interpreted and expressed via corresponding WFFs in $\mathcal{L}$, with no implicit logical self-reference.
Therefore, $S(X,T)$ constitutes an interpretation $E$ of $\mathcal{L}$ over the domain $X\times T$. By (5), both itself and its logical closure are logically consistent. The lemma is thus proved.
Lemma A.1 stipulates that the variables in $\mathcal{L}$ are elements of $X\times T$; thus, in formulas, object and time elements appear in pairs $(x,t)$. In practice, some attributes may be time-invariant; then simplified expressions may be used to describe such attributes without time.
Appendix B Appendix B. Finite Automaton States
Using the formal expression of states, we can study in depth the special and important object of finite automata (Xu, 2024).
Lemma B.1 (State Representation of Finite Automata) For any finite automaton $M$, one can describe its input, output, and state‐transition behavior over a related time set $T$ by the state set $S(M,T)$.
Proof: Let $M=\left\langle Q,R,U,\delta,\lambda\right\rangle$ be a finite automaton, and $T=\left\{t_{i}\middle|i=1,\ldots,n\right\}$ the (strictly increasing) time points at which $M$ performs input, output, or state‐transition actions. By finite‐automaton theory (Editorial Committee of Dictionary of Mathematics, 2002), $Q=\left\{q_{t_{i}}\middle|i=1,\ldots,n-1\right\}$, $R=\left\{r_{t_{i}}\middle|i=2,\ldots,n\right\}$, and $U=\left\{u_{t_{i}}\middle|i=1,\ldots,n\right\}$ are nonempty finite sets of input symbols, output symbols, and states, respectively;
$\delta$: $U\mathbf{\times}Q$ to $U$, is the next‐state function;
$\lambda$: $U\mathbf{\times}Q$ to $R$, is the output function.
For each state $u_{t_{i}}\in U$, define the state predicate ${State}^{3}$ by
$\varphi_{U}(t_{i})={State}^{3}(M,t_{i},u_{t_{i}})$ (B.1)
to assert that at time $t_{i}$, $M$ is in state $u_{t_{i}}$, $i=1,\ldots,n$.
For each input $q_{t_{i}}\in Q$, define the input predicate ${Input}^{3}$ by
$\varphi_{I}(t_{i})={Input}^{3}(M,t_{i},q_{t_{i}})$ (B.2)
to assert that at time $t_{i}$, $M$ receives input $q_{t_{i}}$, $i=1,\ldots,n-1$.
For each output $r_{t_{i}}\in R$, define the output predicate ${Output}^{3}$ by
$\varphi_{O}(t_{i})={Output}^{3}(M,t_{i},r_{t_{i}})$ (B.3)
to assert that at time $t_{i}$, $M$ produces output $r_{t_{i}}$, $i=2,\ldots,n$.
The transition function $\delta$ is captured by the well‐formed formula
$\varphi_{\delta}\left(t_{i}\right)=\varphi_{U}(t_{i})\land\varphi_{I}(t_{i})\rightarrow\exists\delta(u_{t_{i}},q_{t_{i}})\land{Eq}^{2}(u_{t_{i+1}},\delta(u_{t_{i}},q_{t_{i}}))\land\varphi_{U}(t_{i+1})$ (B.4)
where the binary predicate ${Eq}^{2}(x,y)$ asserts $x=y$, $i=1,\ldots,n-1$.
Similarly, the output function $\lambda$ is given by
$\varphi_{\lambda}\left(t_{i}\right)=\varphi_{U}(t_{i})\land\varphi_{I}(t_{i})\rightarrow\exists\lambda(u_{t_{i}},q_{t_{i}})\land{Eq}^{2}(r_{t_{i+1}},\lambda(u_{t_{i}},q_{t_{i}}))\land{IsElementof}^{2}(r_{t_{i+1}},R)\land\varphi_{O}(t_{i+1})$ (B.5)
where ${IsElementof}^{2}(x,X)$ asserts $x$ is an element of $X$, $i=1,\ldots,n-1$.
Thus the state set
$\begin{split}S(M,T)=\{&\varphi_{U}\left(t_{i}\right),\varphi_{I}\left(t_{j}\right),\varphi_{O}\left(t_{k}\right),\varphi_{\delta}\left(t_{j}\right),\varphi_{\lambda}\left(t_{j}\right),\\ &i=1,\ldots,n,j=1,\ldots,n-1,k=2,\ldots,n\}\end{split}$ (B.6)
describes $M$’s inputs, outputs, and state‐transitions over $T$. Here $Cn(X)$ denotes the logical closure of the set of well-formed formulas $X$. This completes the proof.
It follows that every finite automaton can be formalized by the state set given in Theorem 2, which moreover exhibits automatic transition behavior. Hence we introduce:
Definition B.2 (Finite Automaton State) If a state set $S(X,T)$ expresses a finite automaton, then $S(X,T)$ is called a finite automaton state.
Appendix C Appendix C. Model Instance Validation of MLT-MF
@@
C.1 A Simple Feedforward Neural Network Model
We choose a feedforward neural network M with an input layer containing two inputs, a hidden layer containing three neurons, and an output layer containing two outputs (Fig. C.1). Its simplicity helps explain the essence of MLT-MF in a clear and straightforward manner.
Figure 7: Fig. C.1 Schematic diagram of feedforward neural network model M